OpenStudy (anonymous):

If a 30-foot silo casts a 20-foot shadow, what is the approximate measure from the tip of the shadow to the top of the silo? 1,300 feet 36 feet 650 feet 13 feet

OpenStudy (anonymous):



OpenStudy (anonymous):

pythagorean theorem

OpenStudy (anonymous):

:( i fail at this :p

OpenStudy (anonymous):

the square of the 2 legs will equal the longest side we have the 2 legs and we're trying to find the hypotenuse, x

OpenStudy (anonymous):

oh yeah my bad

OpenStudy (anonymous):

so \(30^2+20^2=x^2\)

OpenStudy (anonymous):

1300

OpenStudy (anonymous):

<3 thanx

OpenStudy (anonymous):

not yet bby

OpenStudy (anonymous):

\(30^2+20^2=x^2\\900+400=x^2\\1300=x^2\\x=\sqrt{1300}\)

OpenStudy (anonymous):

you solved for x^2 but we need x

OpenStudy (anonymous):

so divide by 2 or by itself

OpenStudy (anonymous):

take the square root square roots and squaring are inverse operations. one undoes the other

OpenStudy (anonymous):

so just plug in \(\sqrt{1300}\) into a calculator

OpenStudy (anonymous):

so technically 36

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

:) your the best <3

OpenStudy (anonymous):

you're* :p

OpenStudy (anonymous):